A campus club held a bake sale as a fund-raiser, selling coffee, muffins and bacon-and-egg sandwiches. The club members charge led $1 for a cup of coffee, $4 for a muffin and $5 for a bacon-and-egg sandwich. They sold a total of 50 items, raising $133. If the club members sold 2 more muffins than cups of coffee, how many bacon-and-egg sandwiches did they sell?
Answer by stanbon(75887)

A campus club held a bake sale as a fund-raiser, selling coffee, muffins and bacon-and-egg sandwiches. The club members charge led $1 for a cup of coffee, $4 for a muffin and $5 for a bacon-and-egg sandwich. They sold a total of 50 items, raising $133. If the club members sold 2 more muffins than cups of coffee, how many bacon-and-egg sandwiches did they sell?
---------------------
Quantity Eq:: c + m + b = 50 items
Value Eq:::: c + 4m + 5b = 133 dollars
Quantity Eq:: m = c + 2
----------------------------
Since c = m -2 substitute for "c"::
m-2 + m + b = 50
m-2 + 4m+5b = 133
--------------------------
2m + b = 52
5m + 5b = 135
-------
2m + b = 52
m + b = 27
--------
Subtract and solve for "m"
m = 25 # of muffins sold
c = m-2 = 23 (# of cups of coffee sold)
b = 27-m = 4 (# of bacon-eqq-cheese sold)
------------
Cheers,
Stan H.
---------